OpenStudy (honeymonster):

how do i make x the subject of this formula..equation posted in replies..

OpenStudy (honeymonster):

\[3x ^{2} +y -4 =0\]

OpenStudy (anonymous):

isolate x^2 first

OpenStudy (anonymous):

1) add \(4\) 2) subtract \(y\) 3) divide by \(3\) 4) take the square root (don't forget the \(\pm\) )

OpenStudy (honeymonster):

what do i do once i've got here..\[x ^{2}=\frac{ -y+4 }{ 3 }\]..do i square root both sides

hartnn (hartnn):

yes, yes,do it.

OpenStudy (anonymous):

yes, just stick a big fat radical over the whole thing but don't forget the \(\pm\)

OpenStudy (honeymonster):

so thiis would be my answer? \[\sqrt{ x= \frac{ -y+4 }{ 3 }} \]

OpenStudy (honeymonster):

or this?\[\sqrt{\frac{ -y+4 }{ 3 }} = x\]

hartnn (hartnn):

the 2nd one,but u forgot +- sign,even when told twice not to...

OpenStudy (honeymonster):

i dont know what that means..i've had real bad teachers

hartnn (hartnn):

okk,just write it like this \[x=\pm \sqrt{\frac{ -y+4 }{ 3 }} \]

OpenStudy (honeymonster):

oh ok..so i should do that on all square roots like this?

hartnn (hartnn):

yes. because if \[x^2=a\] \[x=\pm a\]

OpenStudy (honeymonster):

ok thanks for your help

hartnn (hartnn):

welcome :)